# Parameter Invariance Analysis of Moment Equations Using Dulmage-Mendelsohn Decomposition

### v1

@@ III-B Dulmage-Mendelsohn decomposition
The truncated stationary moment equations up to the second order moments is then given by
$\begin{bmatrix}0&0&0&-\theta_{4}\\ \theta_{2}&\theta_{5}&-\theta_{5}&0\\ 0&\theta_{3}&0&-\theta_{4}\\ \end{bmatrix}\begin{bmatrix}\mathbb{E}[x_{1}]\\ \mathbb{E}[x_{2}]\\ \mathbb{E}[x_{2}^{2}]\\ \mathbb{E}[x_{1}x_{3}]\end{bmatrix}=-\begin{bmatrix}\theta_{1}\\ 0\\ 0\\ \end{bmatrix}.$ (13)
Here, our goal is to find invariant moments subject to the perturbation to the parameter $\theta_{3}$ of the reaction $R_{3}$. To this end, we formulate the linear equations (7) from eq. (13) and apply the DM decomposition to obtain eq. (12), or
$\begin{bmatrix}\bullet&\bullet&0&0&\bullet&0&0&0\\ 0&0&\bullet&\bullet&0&0&\bullet&0\\ 0&0&0&0&\bullet&\bullet&\bullet&0\\ 0&0&0&0&0&\bullet&0&0\\ 0&0&0&0&0&0&\bullet&\bullet\\ 0&0&0&0&0&0&0&\bullet\end{bmatrix}\begin{bmatrix}\mathbb{E}[x_{2}^{2}]^{(1)}\\ \mathbb{E}[x_{1}]^{(1)}\\ \mathbb{E}[x_{2}^{2}]\\ \mathbb{E}[x_{1}]\\ \mathbb{E}[x_{2}]^{(1)}\\ \mathbb{E}[x_{1}x_{3}]^{(1)}\\ \mathbb{E}[x_{2}]\\ \mathbb{E}[x_{1}x_{3}]\end{bmatrix}=\begin{bmatrix}0\\ 0\\ 0\\ 0\\ 0\\ \bullet\end{bmatrix}.$ (14)
The resulting coefficient matrix becomes upper-triangular, and thus, by applying the backward substitution for linear equations, we can verify that ${\mathbb{E}}[x_{1}x_{3}]^{(1)}=0$ regardless of the parameter value $\bm{\theta}$.
In the next subsection, we consider a graph-based representation of eq. (12) and propose a more systematic procedure to analyze the invariance of moments with respect to parameters.
@@ III-C Graph-based identification of invariant moments
We generalize the idea of the backward substitution shown in Example to systematically identify the invariant moments. To this end, we define $\mathcal{R}_{i}$ and $\mathcal{C}_{i}$ as the set of rows and columns corresponding to the $i$-th diagonal block of the matrix $\Psi$ ($i=0,1,2,\ldots,2\eta$), where the subscript 0 designates the upper-left block $\tilde{A}_{u1}$ of $\Psi$. Corresponding to this partition, we define the $12\eta+$ 1 partitions of $\bm{\nu}_{\Phi}$ and $\bm{b}_{\Psi}$ as
${\bm{\nu}}_{\Psi}=[\tilde{\bm{\nu}}_{0},\tilde{\bm{\nu}}_{1},\dots,\tilde{\bm{\nu}}_{2\eta}]^{\top},\>{\bm{b}}_{\Psi}=[\tilde{\bm{b}}_{0},\tilde{\bm{b}}_{1},\dots,\tilde{\bm{b}}_{2\eta}]^{\top},$ (15)
where $\tilde{\bm{\nu}}_{i+\eta}$ and $\tilde{\bm{\nu}}_{i}\>(i=1,2,\dots,\eta)$ represent the stationary moments and those partial derivative, respectively. In what follows, we interpret these variables as the nodes in a graph $G(\mathcal{V},\mathcal{E})$ with node set $\mathcal{V}$ and edge set $\mathcal{E}$, and we analyze the parameter invariance of the stationary moments through the reachability properteies of the graph. More formally, let $V_{i}$ denote the node corresponding to the variables $\tilde{\bm{\nu}}_{i}$, and we define the set of nodes by $\mathcal{V}:=\{V_{1},V_{2},\ldots,V_{2\eta}\}$. An edge $(V_{q},V_{p})$ is introduced from $V_{q}$ to $V_{p}$ if the corresponding block in $\Psi$ is non-zero, i.e, $\Psi[\mathcal{R}_{p},\mathcal{C}_{q}]\neq O$, where $\Psi[\mathcal{R}_{p},\mathcal{C}_{q}]$ is the $(p,q)$ block of $\Psi$ associated with the row set $\mathcal{R}_{p}$ and the column set $\mathcal{C}_{q}$. Thus, the set of all edges is defined by
:=$\mathcal{E}:=\{(V_{q},V_{p})\in\mathcal{V}\times\mathcal{V}\mid\Psi[\mathcal{R}_{p},\mathcal{C}_{q}]\neq O\}.$ (16)
@@
Consider the graph $G(\mathcal{V},\mathcal{E})$ induced by eq. (12). Define a set of source nodes by $\mathcal{S}:=\{V_{i}\in\mathcal{V}\mid\tilde{\bm{b}}_{i}\neq\bm{0}\}(\subseteq\mathcal{V})$. Then, $\tilde{\bm{\nu}}_{i}=\bm{0}$ holds for every equation (12) if $V_{i}$ is unreachable from all source nodes $\mathcal{S}$. $\tilde{\bm{\nu}}^{(1)}_{r}=\bm{0}$.
@@
In general, the equations for the row set $\mathcal{R}_{i}$ are given by
$\Psi[\mathcal{R}_{i},\mathcal{C}_{i}]\tilde{\bm{\nu}}_{i}+\sum_{q\in\mathcal{N}_{i}}\Psi[\mathcal{R}_{i},\mathcal{C}_{q}]\tilde{\bm{\nu}}_{q}=\tilde{\bm{b}}_{i}$ (17)
where $\mathcal{N}_{i}:=\{j\mid(V_{j},V_{i})\in\mathcal{E}\}$. This implies that $\tilde{\bm{\nu}}_{i}=\bm{0}$ if the variables $\tilde{\bm{\nu}}_{q}\ (q\in\mathcal{N}_{i})$ and $\bm{\tilde{b}}_{i}$ in eq. (17) are zero ()($\star$)( ⋆ ). Since $V_{i}$ is unreachable from the source nodes, $\tilde{\bm{b}}_{q}=0$ for every equation associated with an upstream nodes $V_{q}$ along the directed paths leading to $V_{i}$. Moreover, the equations for the most upstream node $V_{j}$ reduces to $\Psi[\mathcal{R}_{j},\mathcal{C}_{j}]\tilde{\bm{\nu}}_{j}=\bm{0}$, which implies $\tilde{\bm{\nu}}_{j}=\bm{0}$ as $\Psi[\mathcal{R}_{j},\mathcal{C}_{j}]$ is a welldetermined block. Thus, by recursively applying ()($\star$)( ⋆ ) from the most upstream node to $V_{i}$, we have $\tilde{\bm{\nu}}_{i}=0$. ∎
Theorem 1 represents the idea of backward substitution in Example using the concept of unreachability in directed graphs. Since efficient algorithms for determining reachability in directed graphs have been extensively studied [22, 23, 24], it is possible to systematically and efficiently identify invariant stationary moments subject to the perturbation of the parameter $\theta_{k}$. More specifically, $\mathcal{M}_{k}$ can be obtained by applying Theorem 1 to $\tilde{\bm{\nu}}_{i}\>(i=1,2,\dots,\eta)$, by which we can identify the stationary moments whose partial derivatives are structurally determined to be zero.

### v2

@@ III-A Overview of the proposed approach
Figure 1: Identification of invariant moments based on the row and column permutation of eq. (9). The rearranged equation shows $\partial_{3}\mathbb{E}[x_{1}x_{3}]=0$ for all $\bm{\theta}$.
Our goal is to find invariant moments under perturbations in the parameter $\theta_{3}$ of the reaction $R_{3}$. When the moment equations are truncated up to $\sum^{n}_{i=1}\alpha_{i}=1$, eq. (7) is given by
$\!\!\!\!\begin{bmatrix}\!0&\!\!\!0&\!\!\!\!0&\!\!\!\!\!-\theta_{4}&\!\!\!0&\!\!\!0&\!\!\!\!0&\!\!\!\!\!\!0\\ \theta_{2}&\!\!\!\theta_{5}&\!\!\!\!-\theta_{5}&\!\!\!\!\!0&\!\!\!0&\!\!\!0&\!\!\!\!0&\!\!\!\!\!\!0\\ \!0&\!\!\!\theta_{3}&\!\!\!\!0&\!\!\!\!\!-\theta_{4}&\!\!\!0&\!\!\!1&\!\!\!\!0&\!\!\!\!\!\!0\\ \!0&\!\!\!0&\!\!\!\!0&\!\!\!\!\!0&\!\!\!0&\!\!\!0&\!\!\!\!0&\!\!\!\!\!\!-\theta_{4}\\ \!0&\!\!\!0&\!\!\!\!0&\!\!\!\!\!0&\!\!\!\theta_{2}&\!\!\!\theta_{5}&\!\!\!\!-\theta_{5}&\!\!\!\!\!\!0\\ \!0&\!\!\!0&\!\!\!\!0&\!\!\!\!\!0&\!\!\!\!0&\!\!\!\theta_{3}&\!\!\!\!0&\!\!\!\!\!\!-\theta_{4}\\ \end{bmatrix}\!\!\!\begin{bmatrix}{\partial_{3}\mathbb{E}}[x_{1}]\!\!\\ {\partial_{3}\mathbb{E}}[x_{2}]\!\!\\ {\partial_{3}\mathbb{E}}[x^{2}_{2}]\!\!\\ {\partial_{3}\mathbb{E}}[x_{1}x_{3}]\!\\ {\mathbb{E}}[x_{1}]\!\!\\ {\mathbb{E}}[x_{2}]\!\!\\ {\mathbb{E}}[x^{2}_{2}]\!\!\\ {\mathbb{E}}[x_{1}x_{3}]\!\!\end{bmatrix}\!\!=\!\!\begin{bmatrix}0\\ 0\\ 0\\ \!-\theta_{1}\!\\ 0\\ 0\end{bmatrix}\!,\!$ (9)
where $\partial_{3}\mathbb{E}[\cdot]$ represents the derivative of the moment with respect to $\theta_{3}$. This equation is underdetermined, consisting of eight variables and six equations, making it difficult to directly identify the invariant stationary moments by solving the equation.
[…]
A key idea in this example is the rearrangement of eq. (9) into the welldetermined and the underdetermined subequations shown in Fig. 1 and the subsequent backward substitution. In the next subsection, we show that such a transformation can be carried out systematically by using the Dulmage-Mendelsohn (DM) decomposition [18] and that the invariant moments can be found by the reachability analysis of a graph associated with the rearranged equation.
III-B Dulmage-Mendelsohn decomposition
The DM decomposition is a graph theoretic method that provides the finest possible block-triangular decomposition of a matrix, separating its structurally welldetermined part from a given matrix [18, 19]. Specifically, for each parameter value $\bm{\theta}$, we define a bipartite graph $\mathcal{B}(A(\bm{\theta}))$ associated with the matrix $A(\bm{\theta})$, where one set of nodes corresponds to the rows (row nodes) and the other to the columns (column nodes), and an edge connects a row node and a column node if the corresponding entry of $A(\bm{\theta})$ is non-zero. The following lemma follows directly from Theorems 2.1 and 2.2 of [19] with Assumption 2.
Lemma 1
Consider a matrix $A(\bm{\theta})$ satisfying Assumption 2, and the bipartite graph $\mathcal{B}(A(\bm{\theta}))$. Let $M$ be a maximum matching of $\mathcal{B}(A(\bm{\theta}))$. Suppose that some nodes in $\mathcal{B}(A(\bm{\theta}))$ are not reachable from unmatched column nodes via alternating paths with respect to $M$. Then, there exists permutation matrices $P$ and $Q$ such that
$P{A}(\bm{\theta})Q=\begin{bmatrix}{A}_{u}(\bm{\theta})&{A}_{ud}(\bm{\theta})\\ O&{A}_{d}(\bm{\theta})\end{bmatrix},$ (10)
where ${A}_{d}\in\mathbb{R}^{N_{d}\times N_{d}}$ is a block upper-triangular matrix with $\eta$ diagonal blocks satisfying $\mathrm{rank}({A}_{d})=N_{d}$, and ${A}_{u}\in\mathbb{R}^{(M-N_{d})\times(N-N_{d})}$ satisfies $\mathrm{rank}({A}_{u})=M-N_{d}$. Among such transformations, the partition maximizing $\eta$ is unique, and each diagonal block in $A_{d}$ is uniquely determined up to row and column permutations within that block.
The permutation matrices $P$ and $Q$ are obtained by computing the maximum matching of the bipartite graph, followed by a strongly connected component decomposition of the square block $A_{d}$. Polynomial-time algorithms are known for performing these operations [19]. Lemma 1 enables the extraction of the moments and their derivatives that can be fully determined from the underdetermined moment equations when the bipartite graph $\mathcal{B}(A(\bm{\theta}))$ satisfies the assumption in Lemma 1. The next lemma shows that the maximal welldetermined block of $\Phi(\bm{\theta})$ in eq. (7) can be obtained using $A_{d}(\bm{\theta})$.
Lemma 2
Suppose the matrix $A(\bm{\theta})$ satisfies Assumption 2 and admits the decomposition given in eq. (10). There exist row and column permutations that transform $\Phi(\bm{\theta})$ into the upper-triangular block matrix
$\Psi(\bm{\theta})=\begin{bmatrix}{A}_{u1}(\bm{\theta})&{A}_{u2}(\bm{\theta})&{A}_{u3}(\bm{\theta})\\ O&{A}_{d}(\bm{\theta})&{A}^{k}_{d}\\ O&O&{A}_{d}(\bm{\theta})\end{bmatrix},$ (11)
where ${A}_{d}\in\mathbb{R}^{N_{d}\times N_{d}}$ is a block-triangular matrix with $\eta$ diagonal blocks defined in eq. (10), and ${A}_{d}^{k}\in\mathbb{R}^{N_{d}\times N_{d}},{A}_{u1}\in\mathbb{R}^{2(M-N_{d})\times 2(N-N_{d})},{A}_{u2}\in\mathbb{R}^{2(M-N_{d})\times N_{d}}$ and ${A}_{u3}\in\mathbb{R}^{2(M-N_{d})\times N_{d}}$ are real-valued matrices defined accordingly. Moreover, this decomposition maximizes the diagonal blocks among all row and column permutations of $\Phi$, and each diagonal block of $\Psi$ is unique up to row and column permutations within the block.
The proof of this lemma is shown in Appendix -A. Lemma 2 implies that eq. (7) can be systematically rearranged by the DM decomposition to extract the square submatrix corresponding to the welldetermined equations. In the next subsection, we build upon this result by developing a graph-based algorithm that generalizes the backward substitution illustrated in the example in Section III-A. In particular, the proposed algorithm relies solely on the structural information of $\Psi(\bm{\theta})$ and does not require the DM decomposition for each parameter instance, enabling efficient identification of invariant moments for all $\bm{\theta}\in\mathbb{R}^{m}_{+}$.
@@ III-C Graph-based identification of invariant moments
We first define a saturated matrix and vector as follows.
Definition 1
A matrix ${X}(\bm{\theta})$ and a vector $\bm{x}(\bm{\theta})$ parameterized by $\bm{\theta}\in\mathbb{R}_{+}^{m}$ are said to be saturated if its sparsity pattern equals the union of all possible supports, that is, $\mathrm{supp}(X(\bm{\theta}))=\bigcup_{\bm{\theta}_{o}\in\mathbb{R}_{+}^{m}}\mathrm{supp}(X(\bm{\theta}_{o}))$ and $\mathrm{supp}(\bm{x}(\bm{\theta}))=\bigcup_{\bm{\theta}_{o}\in\mathbb{R}_{+}^{m}}\mathrm{supp}(\bm{x}(\bm{\theta}_{o}))$, where $\mathrm{supp}(\cdot)$ is defined in Section I.
It is important to note that the support (sparsity pattern) is identical across all saturated matrices, and likewise for all saturated vectors. In practice, $A(\bm{\theta})$ and $\bm{b}(\bm{\theta})$ in eq. (7) are saturated for most choices of $\bm{\theta}$ as their non-zero patterns are preserved unless the entries vanish due to the cancellation of the parameters. In such typical cases, the sparsity patterns of the saturated matrix $A(\bm{\theta})$ and vector $\bm{b}(\bm{\theta})$ are characterized by constant matrices and vectors as $\bigcup_{j}\mathrm{supp}(A_{j})$ and $\bigcup_{j}\mathrm{supp}(\bm{b}_{j})$, respectively.
Based on this definition, we show a pivotal lemma that leads to the identification of invariant moments based solely on structural information of the equation.
Lemma 3
Suppose the matrix $A(\bm{\theta})$ satisfies Assumption 2. Then, for any $\bm{\theta}\in\mathbb{R}_{+}^{m}$ such that $A(\bm{\theta})$ is saturated and admits the decomposition (10), there exist common permutation matrices $P$ and $Q$ in eq. (10) that transform $\Phi(\bm{\theta})$ into $\Psi(\bm{\theta})$. Moreover, the matrix $\Psi(\bm{\theta})$ obtained in this way is saturated.
Proof:
Since the bipartite graph $\mathcal{B}(A(\bm{\theta}))$ is determined by the support of $A(\bm{\theta})$, the same $P$ and $Q$ apply to eq. (10) for all $\bm{\theta}$ that makes $A(\bm{\theta})$ saturated. By Lemma 2, these $P$ and $Q$ transform $\Phi(\bm{\theta})$ into $\Psi(\bm{\theta})$. The matrix $\Psi(\bm{\theta})$ is saturated since permutations only relabel the rows and columns, and $\Phi(\bm{\theta})$ is saturated when $A(\bm{\theta})$ is saturated. ∎
This lemma implies that a single computation of the DM decomposition for some saturated matrix $A(\bm{\theta})$ suffices to obtain the permutation matrices $P$ and $Q$ leading to the coefficient matrix $\Psi(\bm{\theta})$ of the rearranged equation. More explicitly, we define a linear equation that is obtained by rearranging eq. (7) using the common permutation matrices in Lemma 3 as
$\Psi(\bm{\theta})\bm{\nu}_{\Psi}=\bm{b}_{\Psi}(\bm{\theta}),$ (12)
where $\bm{\nu}_{\Psi}=[{\bm{\nu}}_{0},{\bm{\nu}}_{1},\ldots,{\bm{\nu}}_{2\eta}]^{\top}\in\mathbb{R}^{2N}$ and $\bm{b}_{\Psi}(\bm{\theta})=[{\bm{b}}_{0},{\bm{b}}_{1},\ldots,{\bm{b}}_{2\eta}]^{\top}\in\mathbb{R}^{2M}$ are the vectors obtained by reordering the entries of $\bm{\nu}_{\Phi}$ and $\bm{b}_{\Phi}(\bm{\theta})$, respectively, according to the transformation from the matrix $\Phi(\bm{\theta})$ to the matrix $\Psi(\bm{\theta})$. The sub-vectors ${\bm{\nu}}_{i}$ and $\bm{b}_{i}$ $(i=0,1,\ldots,2\eta)$ are defined in correspondence with the $i$-th diagonal block of $\Psi$, where the subscript 0 refers to the upper-left block ${A}_{u1}$ of $\Psi$. By construction, ${\bm{\nu}}_{i+\eta}$ and ${\bm{\nu}}_{i}\>(i=1,2,\dots,\eta)$ represent the stationary moments and their partial derivatives, respectively.
In what follows, we show that the identification of invariant moments reduces to a reachability analysis on a graph induced by the structure of eq. (12). To this end, we define a graph $\mathcal{G}(\mathcal{V},\mathcal{E})$ with node set $\mathcal{V}:=\{V_{1},V_{2},\ldots,V_{2\eta}\}$ and edge set $\mathcal{E}$ that encodes the structure of the sparsity pattern of $\Psi$, where each node $V_{i}\in\mathcal{V}$ corresponds to each variable ${\bm{\nu}}_{i}$, and an edge $(V_{q},V_{p})$ is introduced from $V_{q}$ to $V_{p}$ if the corresponding block in $\Psi$ is non-zero. Specifically, the edge set is defined by
:=$\mathcal{E}:=\{(V_{q},V_{p})\in\mathcal{V}\times\mathcal{V}\mid\Psi[\mathcal{R}_{p},\mathcal{C}_{q}]\neq O\},$ (13)
where $\mathcal{R}_{i}$ and $\mathcal{C}_{i}$ represent the row and column indices corresponding to the $i$-th diagonal block of the matrix $\Psi$ $(i=0,1,\ldots,2\eta)$, and $\Psi[\mathcal{R}_{p},\mathcal{C}_{q}]$ is the $(p,q)$ block of $\Psi$ specified by the row and column indices $\mathcal{R}_{p}$ and $\mathcal{C}_{q}$. In general, the topology of the graph $\mathcal{G}(\mathcal{V},\mathcal{E})$ is uniquely determined for all $\bm{\theta}$ such that $A(\bm{\theta})$ is saturated since $\Psi(\bm{\theta})$ is also saturated due to Lemma 3, and the sparsity pattern of saturated matrices are identical.
@@
Suppose Assumption 2 holds. Let $\bm{\theta}^{*}$ be the parameter such that $A(\bm{\theta}^{*})$ admits the decomposition (10), and $A(\bm{\theta}^{*})$ and $\bm{b}(\bm{\theta}^{*})$ are saturated. Define the graph $\mathcal{G}(\mathcal{V},\mathcal{E})$ induced by eq. (12) with $\bm{\theta}^{*}$, and let the set of source nodes by $\mathcal{S}:=\{V_{i}\in\mathcal{V}\mid{\bm{b}}_{i}\neq\bm{0}\}(\subseteq\mathcal{V})$. Then, ${\bm{\nu}}_{i}=\bm{0}$ in eq. (12) holds for any $\bm{\theta}\in\mathbb{R}^{m}_{+}$ if $V_{i}$ is unreachable from all source nodes $\mathcal{S}$.
@@
The proof is divided in two parts. (i) For a fixed $\bm{\theta}^{*}$, consider eq. (12). The equation corresponding to the row set $\mathcal{R}_{i}$ is
$\Psi[\mathcal{R}_{i},\mathcal{C}_{i}]{\bm{\nu}}_{i}+\sum_{q\in\mathcal{N}_{i}}\Psi[\mathcal{R}_{i},\mathcal{C}_{q}]{\bm{\nu}}_{q}={\bm{b}}_{i}$ (14)
where $\mathcal{N}_{i}:=\{j\mid(V_{j},V_{i})\in\mathcal{E}\}$. It follows that ${\bm{\nu}}_{i}=\bm{0}$ if ${\bm{\nu}}_{q}$ for $q\in\mathcal{N}_{i}$ and $\bm{{b}}_{i}$ are zero ()($\star$)( ⋆ ). Since $V_{i}$ is unreachable from the source nodes, ${\bm{b}}_{q}=0$ for every equation associated with upstream nodes $V_{q}$ along the directed paths leading to $V_{i}$. Moreover, the equations for the most upstream node $V_{j}$ reduces to $\Psi[\mathcal{R}_{j},\mathcal{C}_{j}]{\bm{\nu}}_{j}=\bm{0}$, which implies ${\bm{\nu}}_{j}=\bm{0}$ as $\Psi[\mathcal{R}_{j},\mathcal{C}_{j}]$ is a full rank square matrix due to Assumption 2. Thus, by recursively applying ()($\star$)( ⋆ ) from the most upstream node to $V_{i}$, we have ${\bm{\nu}}_{i}=0$.
(ii) Next, we show that the same result holds for any $\bm{\theta}\in\mathbb{R}_{+}^{m}$. Consider $\bm{\theta}$ such that $A(\bm{\theta})$ and $\bm{b}(\bm{\theta})$ are saturated. Then, the graph and the source nodes induced by $\Psi(\bm{\theta})$ and $\bm{b}(\bm{\theta})$ are identical to those induced by $\Psi(\bm{\theta}^{*})$ and $\bm{b}(\bm{\theta}^{*})$ since $\Psi(\bm{\theta})$ is saturated by Lemma 3, and saturated matrices and vectors have the same sparsity pattern regardless of the parameter $\bm{\theta}$. Thus, by the same argument as in (i), ${\bm{\nu}}_{i}=\bm{0}$ holds. For other $\bm{\theta}$, where either $A(\bm{\theta})$ or $\bm{b}(\theta)$ is not saturated, some entries of $\Psi(\bm{\theta})$ or $\bm{b}(\theta)$ that were not zero in the saturated case become zero. As a result, the associated graph and the source nodes are obtained by removing the corresponding edges and source nodes from the graph $\mathcal{G}(\mathcal{V},\mathcal{E})$ for the saturated case. As these are subsets of those in the saturated case, unreachability is preserved. Thus, the same argument as in (i) leads to ${\bm{\nu}}_{i}=\bm{0}$. ∎
Theorem 1 shows that, based on the idea of backward substitution, the analysis of parameter invariance of stationary moments can be reduced to a reachability problem in a directed graph. The Floyd-Warshall algorithm provides an efficient way to compute the reachability from multiple source nodes $\mathcal{S}$ in $O(|\mathcal{V}|^{3})$ time with respect to the number of nodes $|\mathcal{V}|$[22].
Remark: The proposed method does not guarantee identification of all invariant moments $\mathcal{M}_{k}$. This limitation stems from two main factors: (i) some invariant moments may not be identified if they depend on higher order terms omitted by the truncation of moment equation, (ii) invariance arising from dependencies among matrix entries cannot be detected as the proposed approach uses only the sparsity pattern.